# TRACE: A Framework for Analyzing and Enhancing Stepwise Reasoning in Vision-Language Models

### v1

@@ 4.3.4 First Failure Step (FFS) Analysis
In Table 2, the model predicts $\angle AOC=130^{\circ}$ while the correct majority answer is $80^{\circ}$, marking the First Failure Step (FFS). This demonstrates that even a single incorrect intermediate answer can propagate to the final prediction ($\angle A=65^{\circ}$ vs. correct $25^{\circ}$). By including such pivotal sub-questions in the ARS, TRACE can reliably localize reasoning failures, highlight critical dependencies, and provide actionable insights for debugging and improving VLM performance on multi-step tasks. Appendix LABEL:ffs-exampleIII presents further examples of inconsistent reasoning patterns, extending beyond simple deviations from the majority.
@@ 6.3 ARS Quality
After filtering ARS sets for quality and leakage, key statistics for exploration and exploitation paths are reported in Table 8 and 7. On average, MMMUPro requires about six sub-questions per problem, compared to roughly four for the TIGER dataset. This suggests that MMMUPro problems demand more reasoning steps. The higher complexity of MMMUPro is also reflected in accuracy: 54.3% versus 79.9% for TIGER (Table 6). Thus, ARS statistics not only capture structural properties of reasoning paths but also provide insights into dataset difficulty. Consistent with this, the average sub-question dependency is higher for MMMUPro than for TIGER, further indicating its greater reasoning complexity. Figures 7 and 6 present the distributions of sub-question dependencies for the TIGER and MMMUPro datasets, respectively. The dependency graph illustrates that the initial three questions function as the most frequent dependency nodes for all subsequent sub-questions, confirming the expected hierarchical structure of the task. This centralized structure highlights the sensitivity of the overall reasoning chain to the model’s performance on these foundational steps, as errors introduced here are likely to propagate to dependent sub-questions.

### v2

@@ 4.3.4 First Failure Step (FFS) Analysis
In Table 2, the model predicts $\angle AOC=130^{\circ}$ while the correct majority answer is $80^{\circ}$, marking the First Failure Step (FFS). This demonstrates that even a single incorrect intermediate answer can propagate to the final prediction ($\angle A=65^{\circ}$ vs. correct $25^{\circ}$). By including such pivotal sub-questions in the ARS, TRACE can reliably localize reasoning failures, highlight critical dependencies, and provide actionable insights for debugging and improving VLM performance on multi-step tasks. Appendix 5 presents further examples of inconsistent reasoning patterns, extending beyond simple deviations from the majority.
@@ 5 Conclusion and Future Work
FFS Example II
Consider another case where the model deviates from the majority at an early sub-question. Here, the final answer is incorrect, with the first error occurring at Q2, highlighting how early mistakes propagate and why identifying FFS is crucial for debugging multi-step reasoning.
Main Question:
In the Cartesian coordinate system xOyxOy, a circle centered at the origin $O$ passes through the point $A(13,0)$. The line y=kx3k+4y=kx-3k+4 intersects the circle $\odot O$ at points $B$ and $C$. What is the minimum length of the chord BCBC?
ARS Sub-question + Main Question Path Answer / Majority Answer
Q1: What is the radius of the circle centered at $O$? 13 / 13
Q2: What is the distance from $O$ to the line y=kx3k+4y=kx-3k+4? 4 / 5 (First Incorrect)
Main Question: What is the minimum length of chord BCBC? $6\sqrt{17}$ / 24
Table 3: Sub-questions and corresponding path answers vs. majority answers for the second FFS example.
Why the Distance to the Line is Critical The correct computation of the distance from $O$ to the line y=kx3k+4y=kx-3k+4 (Q2) is essential, as it directly determines the chord length BCBC. Geometric Relationship: The perpendicular distance $d$ from $O$ to the line sets the position of the chord within the circle. For a circle of radius $r$, the chord length is: $|BC|=2\sqrt{r^{2}-d^{2}}$ With $r=13$ and $d=5$, the minimum chord length is: $|BC|=2\sqrt{13^{2}-5^{2}}=2\sqrt{169-25}=24$
Illustrative Example III
Consider a reasoning path from the MMMUPro dataset where the final answer is incorrect. FFS identifies the first sub-question in the path where the model’s response deviates from the expected majority. In this example, the GPT-4.1 model provides inconsistent answers to Q3 across iterations—once identifying $B$ and $C$ are the foci” and another time AA and $C$. This inconsistency marks Q3 as the initial point of reasoning failure. This illustrates how ARS dependencies help isolate faulty visual reasoning steps.
Main Question:
The elliptical orbit of a planet around the Sun is shown in the diagram. Which of the following statements is true?
ARS Sub-question + Main Question Path Answer I/ Path Answer II
Q1: What is the shape of the planet’s orbit as shown in the diagram? ellipse / ellipse
Q2: How many focal points does the elliptical orbit have in the diagram? 2 / 2
Q3: Which labeled points (A, B, C, D, E) are located at the foci of the ellipse? B and C / A and C (First Incorrect)
Q4: Is the Sun located at one of the foci of the elliptical orbit according to the diagram? Yes / Yes
Q5: Is the eccentricity of the orbit equal to zero? No / No
Q6: Is the eccentricity of the orbit equal to one? No / No
Q7: Is the eccentricity of the orbit greater than one? No / No
Q8: Is the eccentricity of the orbit less than zero? No / No
Main Question: Which of the following statements is true? the sun might be at point C / the sun might be at point A or C (both incorrect)
Table 4: Sub-questions and corresponding path answers vs. majority answers for a MMMUPro orbit example. The first reasoning failure occurs at Q3.
Why Identifying the Foci Matters The sub-question “Which labeled points are located at the foci of the ellipse?” is pivotal for determining whether the Sun is correctly placed in the diagram. Geometric Relationship: An ellipse has exactly two foci. Kepler’s First Law: A planet orbits the Sun in an ellipse with the Sun at one focus. Misidentifying the foci propagates to an incorrect placement of the Sun and therefore an incorrect final answer.
Illustrative Example IV
This example presents a First Failure Step (FFS) in a geometric ratio computation. The model’s reasoning fails at Q3, which propagates to an incorrect final area for $\triangle ABC$.
Main Question:
In $\triangle ABC$, $E$ is the midpoint of ACAC, and $D$ lies on BCBC such that $BD:CD=2:3$. Lines ADAD and BEBE intersect at $O$. If $S_{\triangle AOE}-S_{\triangle BOD}=1$, find the area of $\triangle ABC$.
ARS Sub-question + Main Question Path Answer / Majority Answer
Q1: What is the ratio of BDBD to CDCD? $(2/3)$ / $(2/3)$
Q2: Is $E$ the midpoint of ACAC? Yes / Yes
Q3: What is the ratio of areas $\triangle AOE:\triangle BOD$? 4 / $(5/4)$ (First Incorrect)
Q5: What is the area of $\triangle ABE$? 6 / 5
Q6: What is the area of $\triangle ABD$? $(56/9)$ / 4
Main Question: Area of $\triangle ABC$? $(140/9)$ / 10
Table 5: Sub-questions and path vs. majority answers
First Failure Step (FFS):
The model first fails at Q3, incorrectly computing the ratio of areas $\triangle AOE:\triangle BOD$ as 44 instead of $5/4$. This miscalculation propagates to subsequent area computations, leading to an incorrect area for $\triangle ABC$.
Why Area Ratios Determine the Final Answer The ratio $\dfrac{S_{\triangle AOE}}{S_{\triangle BOD}}$ links the condition $S_{\triangle AOE}-S_{\triangle BOD}=1$ to the absolute area of $\triangle ABC$. Geometric Relationship: Intersection of cevians ADAD and BEBE defines proportional sub-triangles. Correct ratio ($5/4$) allows solving for $S_{\triangle AOE}$ and $S_{\triangle BOD}$. Summing sub-triangle areas scales to $S_{\triangle ABC}=10$. ARS Dependency Justification: Q3 directly governs the chain of reasoning. An incorrect ratio leads to errors in all subsequent area computations.
Illustrative Example V
This example illustrates a First Failure Step (FFS) in coordinate reasoning. The model first fails at Q5, leading to an incorrect final reflection point $B^{\prime}$.
Main Question:
In rectangle OABCOABC, OAOA and OCOC lie on the $x$- and $y$-axes. $B=(3,2)$, points $D$ and $E$ lie on ABAB and BCBC with $BD=BE=1$. Folding $\triangle BDE$ along DEDE maps $B$ to $B^{\prime}$. Find the coordinates of $B^{\prime}$.
ARS Sub-question + Main Question Path Answer / Majority Answer
Q1: Coordinates of $B$? $(3,2)$ / $(3,2)$
Q2: Length of BDBD? 11 / 1
Q3: Length of BEBE? 11 / 1
Q4: Coordinates of $D$? $(3,1)$ / $(3,1)$
Q5: Coordinates of $E$? $(3,1)$ / $(2,2)$ (First Incorrect)
Q6: Slope of DEDE? Undefined / 1-1
Q7: Equation of line DEDE? $x=3$ / y=x+4y=-x+4
Main Question: Coordinates of $B^{\prime}$? $(3,0)$ / $(2,1)$
Table 6: Sub-questions and path vs. majority answers.
First Failure Step (FFS):
The model first fails at Q5 by incorrectly placing $E$ at $(3,1)$ instead of $(2,2)$. This error propagates to Q6 (slope), Q7 (line equation), and the final reflection, producing an incorrect $B^{\prime}$. Identifying Q5 as the FFS reveals the cascading effect of a single misstep in coordinate placement.
Why Correct Placement of $E$ is Critical Point $E$ lies on BCBC, one unit from $B(3,2)$. Correct placement gives $E=(2,2)$. Consequences: Line DEDE has slope 1-1 and equation y=x+4y=-x+4. Folding across DEDE maps $B(3,2)$ to $B^{\prime}=(2,1)$. An incorrect coordinate at Q5 misaligns all subsequent computations. ARS Dependency Justification: Q5 directly influences the line DEDE and the final reflection. Local errors propagate through the reasoning chain, demonstrating how the FFS captures the earliest point of failure.
@@ 6.3 ARS Quality
After filtering ARS sets for quality and leakage, key statistics for exploration and exploitation paths are reported in Table 12 and 11. On average, MMMUPro requires about six sub-questions per problem, compared to roughly four for the TIGER dataset. This suggests that MMMUPro problems demand more reasoning steps. The higher complexity of MMMUPro is also reflected in accuracy: 54.3% versus 79.9% for TIGER (Table 10). Thus, ARS statistics not only capture structural properties of reasoning paths but also provide insights into dataset difficulty. Consistent with this, the average sub-question dependency is higher for MMMUPro than for TIGER, further indicating its greater reasoning complexity. Figures 7 and 6 present the distributions of sub-question dependencies for the TIGER and MMMUPro datasets, respectively. The dependency graph illustrates that the initial three questions function as the most frequent dependency nodes for all subsequent sub-questions, confirming the expected hierarchical structure of the task. This centralized structure highlights the sensitivity of the overall reasoning chain to the model’s performance on these foundational steps, as errors introduced here are likely to propagate to dependent sub-questions.